OpenStudy (anonymous):

I include the lession example, for the small angle I can do no problem , but for big number I can solve In circle C, point T is randomly selected so that it does not coincide with point P, segment SQ is a diameter of circle C. 1)Find the probability that m

OpenStudy (anonymous):

look like limiting word,I update more detail . Sin <ABD must intercept arc AB the probability that m<ADB =30 is the same as the probability of D is located on arc ADB is 300/360 or 5/6 . So the probability that m<ADB =30 is also 5/6 I do 360-90=270 270/360=3/4 I don't know how get answer is:1